OpenStudy (anonymous):

Can you simplify this: sqrt((9t^4)+(18t^2)+9)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Could you show me how?

OpenStudy (anonymous):

factor out a 9 first and get \[\sqrt{9(t^4+2t^2+1)}\]

OpenStudy (anonymous):

then note that \(t^4+2t^2+1\) is itself a perfect square, namely \[t^4+2t^2+1=(t^2+1)^2\]

OpenStudy (anonymous):

this gives you \[\sqrt{9(t^2+1)^2}\]

OpenStudy (anonymous):

then take the square root and get \(3(t^2+1)\)